the length of viollas bathroom is 2 feet longer than twice the width. if the diagonal measures 13 feet then what are the length and width. ive tried to solve this so many different ways but i cant get the right answer.
Answer by scott8148(6628)

let x=width , then 2x+2=length

by Pythagoras ___ x^2 + (2x+2)^2 = 13^2 ___ x^2 + 4x^2 + 8x + 4 = 169

5x^2 + 8x - 165 = 0

factoring ___ (5x + 33)(x - 5) = 0 ___ negative solution is not realistic

x - 5 = 0 ___ x = 5